On a sheet of paper, graph a line segment from starting point S(2,5) to ending point E(-1,-3) and then perform the following:
a. What will be the new starting and ending points if you reflected this line segment through the line y = x ?
b. Find the point on this new (reflected) segment which is 3/4 of the way from the starting point to the ending point. What quadrant is this point in?
c. Explain how you arrived at your answers.

Answer by greenestamps(13244)



a. Reflecting the graph (ANY graph) in the line y=x simply switches the x and y values. So the new starting point is (5,2) and the new ending point is (-3,-1).

b. To find the point on the new line 3/4 of the way from the starting point to the ending point, compute the x and y values of the point sepearately.

From starting point to ending point, the change in x is -8 (from 5 to -3); 3/4 of -8 is -6; -6 from the starting x value of 5 is 5-6 = -1. The x coordinate of the point we are looking for is -1.

From the starting point to the ending point, the change in y is -3 (from 2 to -1); 3/4 of -3 is -9/4; -9/4 from the starting y value of 2 is 2-9/4 = 8/4-9/4 = -1/4. The y coordinate of the point is -1/4.

So the point is (-1,-1/4); that is in quadrant III.

c. The explanations are given above for each part.